OpenStudy (anonymous):

can you explain in words for me to understand, additive identity, additive inverse, associative property, commutative property, distributive property, multiplicative identity, and multiplicative inverse.

OpenStudy (anonymous):

commutative property of add ition refers to how you can swop over the numbers,letters eg a + b = b + a commutative of multiplication eg ab = ba

OpenStudy (anonymous):

distributive property is illusstrated by x ( 2 + y) = 2x + xy

OpenStudy (anonymous):

Are u trying to establish the axiomatic structure applicable to numbers or are u more interested in group theory generally. PROPERLY understanding all those things is more than a post or two in here, I'm afraid.

OpenStudy (anonymous):

i am just looking for a basic explination of all of those for my notes. My teacher gave us long definitions that i can hardly understand. jimmyrep; Thankyouu!:)

OpenStudy (anonymous):

Thankyouu!